OpenStudy (anonymous):

Solve the system of linear equations below. 4x - 3y = 17 5x + 4y = 60

OpenStudy (anonymous):

what happens when you add those two equations?

OpenStudy (anonymous):

yu tell me

OpenStudy (anonymous):

4x - 3y = 17 ---> Equation 1 5x + 4y = 60 ---> Equation 2 4(equation 1) + 3(equation 2) 4(4x-3y) = 4(17) +3(5x+4y)=3(60) ----------------------------------- (16+15)x + (-12 +12)y = (17)(4)+3(60) 31 x + 0 = 68 + 180 31 x = 248 x = 248 /31 = 8 Now subsitute x= 8 in either equation 1 or 2 to get your y value 4(8) - 3 y =17 32 - 3y =17 32-17 = + 3y y = 15/3 = 5